TA= Total Acidity

To determine TA, a 15 g portion of a juice was taken, and

17.5 mL of 0.085 N NaOHwas used to titrate the total acid.

What is the TA of the juice, expressed as % wt of citric acid?

(Citric acid MW = 192, with 3 dissociable H+)
Answer #1

total moles of H+ ions reacted = 0.085 * 17.5 = 1.487*10^-3 moles

since 1 mole of citric acid contain 3 moles of H+ ions

so moles of citric acid reacted = 1.487 * 10^-3 / 3 = 0.496 *10^-3 moles

amount of citric acid reacted = 0.496*10^-3 * 192 = 95.2 *10^-3 g

% citric acid = (amount of citric acid reacted / amount taken citric acid ) * 100

= (95.2*10^-3 / 15 )*100

= 0.635 %